OpenStudy (anonymous):

47. Multiply and simplify if possible. (problem below)

OpenStudy (anonymous):

\[\sqrt{7x}(\sqrt{x}-7\sqrt{7})\]

OpenStudy (anonymous):

@AravindG

OpenStudy (zehanz):

Use the distributive property: a(b-c)=ab-ac, so:\[\sqrt{7x}\sqrt{x}-\sqrt{7x}7\sqrt{7}=...\]

OpenStudy (anonymous):

idk how to solve it i got a decimal when i used the calculator.

OpenStudy (zehanz):

Just multiply as follows: √a√b=√(ab), so: √(7x)√x=√(7x²) and √(7x)7√7=7√(49x) So you now have :√(7x²) - 7√(49x) Both these terms can be simplified a little more, but do not use the calculator!

OpenStudy (anonymous):

so would \[\sqrt{7x^3}=\sqrt{49}\] be the answer???

OpenStudy (anonymous):

xsqrt7-7sqrtx is the answer right?

OpenStudy (zehanz):

No, you cannot subtract these different terms. Your last answer is almost right! √(7x²)=√7√x²=x√7 7√(49x)=7√49√x=7*7√x=49√x Final answer: x√7-49√x